A ball rebounds one-half the height from which it is dropped. The ball is dropped from a height of 160 ft. and keeps on bouncing. What is the total vertical distance the ball will travel from the moment it is dropped to the moment it hits the floor for the fifth time?
I have no clue how to solve this.

Answer by josmiceli(19441)

The ball falls 160, bouces once, comes back up to half-way, so far
it's travelled 160 + 80 = 240.
Now it drops from 80, bounces for the 2nd time, and comes back up to half
of 80, or 40. So far it's travelled 240 + 80 + 40 = 360.
It drops from 40, bounces the 3rd time, comes back up to 20.
So far, it's travelled 360 + 40 + 20 = 420.
Now it drops from 20, bounces the 4th time, comes back up to 10
So far, it's travelled 420 + 20 + 10 = 450
Now it drops from 10 and hits the floor for the 5th time.
The total distance is 450 + 10 = 460 feet
What we just did was add up
160
80
80
40
40
20
20
10
10
-----
460